Can a kind tutor please tell me if I did these difficult homework problems correctly?
7. Of the 100 people who were given a new sleeping aid, 30 experienced a light headache, 18 experienced a reduced appetite, and 60 experienced neither side-effect. How many people experienced both a light headache and a reduced appetite?
60-48 - 12?
8. A quiz consists of 6 multiple-choice questions, each has three possible answers. In how many different ways can the quiz be answered if one answer is chosen for each question?
3^6= 729?
9. A code consists of two letters chosen from the English alphabet, followed by 4 digits. If the first letter must be a K or an M, the second letter can be any one of the 26 letters, how many different codes are possible?
2 x 26= 52 x 10^4= 52,000?
10.In how many ways can two math books and three English books be arranged in a row if the two math books must be placed next to each other?
4!?

Answer by Edwin McCravy(20086)

Can a kind tutor please tell me if I did these difficult homework problems correctly?
7. Of the 100 people who were given a new sleeping aid,
30 experienced a light headache, 18 experienced a reduced
appetite, and 60 experienced neither side-effect. How many
people experienced both a light headache and a reduced appetite?
60-48 - 12?
No that's wrong

Let A = set of those with a light headache
Let B = set of those with a reduced appetite

N(A OR B) = N(A) + N(B) - N(A AND B)

Since 60 experienced neither, 100-60 or 40 experienced
one OR the other.

So N(A OR B) = 40

We are given N(A) = 30 and N(B) = 18 
We want to find N(A AND B).  Let that = x:

N(A OR B) = N(A) + N(B) - N(A AND B)
       40 =  30  +  18  -    x

       40 =  48 - x

        x = 8 

-------------------------------------------------         

8. A quiz consists of 6 multiple-choice questions,
each has three possible answers. In how many different
ways can the quiz be answered if one answer is chosen
for each question?
3^6= 729?
That's correct.

-------------------------------------------------

9. A code consists of two letters chosen from the English
alphabet, followed by 4 digits. If the first letter must
be a K or an M, the second letter can be any one of the
26 letters, how many different codes are possible?
2 x 26= 52 x 10^4= 52,000?
You shouldn't put the intermediate step of 52 inside
the equation.  You should have 

2 x 26 x 10^4 = 

and then the answer is 520,000, not what you have, You
need another 0.

-------------------------------------------------

10.In how many ways can two math books and three
English books be arranged in a row if the two math
books must be placed next to each other?
4!?
No, that's wrong.

We have one of 4 configurations:  

MMEEE, EMMEE, EEMME, EEEMM

That's 4 choices for the configurations

For each of those configurations we can arrange the
2 math books 2! ways. That's 4 x 2! ways to choose 
the configuration and the math books.

For each of those 4 x 2! ways, there are 3! ways 
to arrange the 3 English books.

So the answer is 4 x 2! x 3! = 4 x 2 x 6 = 48 ways.

Edwin